### v4

# MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning

Abstract
Deriving meaningful representations from complex, high-dimensional data in unsupervised settings is crucial across diverse machine learning applications. This paper introduces a framework for multi-scale graph network embedding based on spectral graph wavelets that employs a contrastive learning approach. We theoretically show that in Paley-Wiener spaces on combinatorial graphs, the spectral graph wavelets operator provides greater flexibility and control over smoothness compared to the Laplacian operator, motivating our approach. An additional key advantage of the proposed embedding is its ability to establish a correspondence between the embedding and input feature spaces, enabling the derivation of feature importance. We validate the effectiveness of our graph embedding framework on multiple public datasets across various downstream tasks, including clustering and unsupervised feature importance.
1 Introduction
Graph Embeddings and Manifold Learning [29, 3, 34, 35, 5] are pivotal in analyzing complex data structures prevalent in diverse machine learning applications. The emergent representations from these techniques offer insights into the underlying data structure in conjunction with downstream tasks like clustering and visualization and are particularly valuable when labels are unavailable or unreliable. However, limitations of these techniques include reliance on low frequencies of the graph Laplacian and $k$-Nearest Neighbors graph connectivity, which provide limited information, including in recent contrastive learning methods such as Uniform Manifold Approximation and Projection (UMAP). To address these limitations, this paper proposes three main contributions.
(1) A framework employing multi-scale graph representation using 3D based tensor and optimization using contrastive learning which enhances the expressiveness of an embedding by optimizing low and high-frequency information, capturing intricate data details. (2) An application of manifold interpretation to feature importance allows each dimension of our proposed embedding to be uniquely traced back and associated with the original features. (3) A characterization of the theoretical properties of the representation power of the spectral graph wavelets (SGW) operator by considering functions sampled from the Paley-Wiener spaces [27] on combinatorial graphs and a demonstration of the flexibility and better control over smoothness properties of the SGW operator in comparison to the Laplacian operator.
A key distinction between our approach and many manifold learning methods such as UMAP is that we leverage multi-scale graph representation (using SGW) to increase the dimensionality of the data thus enhancing data representation learning. We use stochastic gradient descent (SGD) to optimize the embedding space and employ a novel contrastive learning approach using a 3D-tensor to capture both low and high frequencies.
Generally, manifold learning methods perform nonlinear dimensionality reduction by embedding the data into lower-dimensional spaces, although some approaches may embed the data into higher dimensions [16, 31]. A common limitation of these and other graph embedding methods is their lack of direct connections to the input features, which restricts their ability to produce interpretable results. This can be crucial in fields like finance, where behavioral analysis must be summarized using key features to guide decisions, or in biology, where interpretable features are essential for understanding complex systems. Our proposed approach establishes a mapping between the original features and the constructed optimized embeddings and enables the measurement of the importance of the original features with respect to the embedding space. To the best of our knowledge, current manifold learning techniques produce embeddings lacking explicit interpretation. In contrast, our approach demonstrates the ability to generate interpretable embedding representations, while also being competitive with the state-of-the-art graph embeddings across applications such as finance, vision, and biology.
2 Related Work
Extensive efforts have been dedicated to exploring non-linear dimensionality reduction aimed at achieving a low-dimensional embedding that preserves the underlying manifold structure. Early manifold learning techniques include Locally Linear Embedding (LLE) [29], Laplacian Eigenmaps [3], Isomap [34], Diffusion Maps [4], and t-distributed Stochastic Neighbor Embedding (t-SNE) [35].
Balancing local and global structure in the resulting embeddings is challenging in the presence of noise. To address the challenge, researchers have introduced de-noising methods [20, 12, 8, 11, 7, 13, 9, 10]. These techniques seek to enhance algorithms robustness, thereby broadening the applicability of these techniques in real-world scenarios.
UMAP [25] has proven effective in providing vivid visualizations and cluster separation, yet its reliance on Laplacian Eigenmaps initialization and negative sampling optimization has limitations. Laplacian Eigenmaps initialization focuses on low-frequency signal overlooking higher-frequency patterns. It uniformly samples edges without considering topological importance, overemphasizing graph connectivity and introducing further distortion.
Recent work [6] further highlights that negative sampling has a significant impact on the effective loss and introduces distortions with t-SNE and UMAP which motivates the exploration of alternate sampling schemes. Other relevant studies include [38] which investigates competing negative sampling strategies, [32] on generalization error bounds , and [23] on topological and sampling bias due to high degree nodes.
The lack of explicit mapping linking the original high dimensional dataset to its low dimensional embedding has prevented their use in areas where interpretability is critical. Establishing this link clearly would lead to a better understanding of the original feature space and improve feature selection, understanding of the resulting visualizations, and robustness to noise [21, 15, 39].
SGW serve as an efficient tool for localization in both the spectral and vertex domains, playing a crucial role in our method (see Section 3.1 for details). SGW are constructed by applying a set of bandpass filters to the eigenvalues of the graph Laplacian, enabling multi-resolution representation of graph signals. Also related to our approach are graph embedding and multi-scale framework using Diffusion Wavelets [28] that embed the data in higher dimensions such as the Geometric Scattering for Graph Data Analysis [17] and the Graph Scattering Transform [16] that extract graph features, which are then employed in graph classification tasks.
2.1 Preliminaries
Consider a set of points $\mathbf{x}=\left\{\mathbf{x}_{i}\right\},\,i=1,...N,\mathbf{x}_{i}\in\mathbb{R}^{D}$, which are sampled from an unknown manifold $M\subset\mathbb{R}^{D}$. An undirected, weighted graph $G=(V,\mathbf{W})$, is constructed over $\mathbf{x}$, where $V$ corresponds to the nodes and $\mathbf{W}$ to the set of edges on the graph. The adjacency matrix $\mathbf{W}=(w_{ij})$ consists of the weights $w_{ij}$ between node $i$ and node $j$. The weights can be chosen using different techniques, such as the Gaussian kernel function between the points $\mathbf{x}_{i},\mathbf{x}_{j}$ using the $k$ nearest neighbors of $\mathbf{x}_{i}$, kNN$(\mathbf{x}_{i})$ (which denotes the $k$ nearest neighbors) or adaptive graph construction as proposed in [25]. $\mathbf{L}$ denotes the combinatorial graph Laplacian, defined as $\mathbf{L}=\mathbf{D}-\mathbf{W}$, with $\mathbf{D}$ the diagonal degree matrix with entries $d_{ii}=d(i)$. The degree $d(i)$ of vertex $i$ is defined as the sum of weights of edges that are connected to $i$. The normalized Laplacian is defined as ${\mathbf{L}}_{N}=\mathbf{D}^{-\frac{1}{2}}\mathbf{L}\mathbf{D}^{-\frac{1}{2}}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{W}\mathbf{D}^{-\frac{1}{2}}$ and its real eigenvalues are in the interval $[0,2]$. The eigenvalues and eigenvectors of $\mathbf{L}$ are $\lambda_{1},\ldots,\lambda_{N}$ and $\mathbf{\phi}_{1},\ldots,\phi_{N}$, respectively. In this work we use the normalized Laplacian $\mathbf{L}_{N}$, which from now on we refer to as $\mathbf{L}$ for simplicity. We will denote the matrix representations of the eigenvectors and eigenvalues of the graph Lapacian $\mathbf{L}$ as $\Phi,\lambda$, respectively. Our approach assumes that all graph signals $f_{r}\in\mathbb{R}^{N}$ used to create the embedding correspond to the input coordinates, $\mathbf{x}=\left(f_{1},f_{2},..,f_{r},..f_{D}\right)$. The output embedding space provides a multi-scale representation approximating the intrinsic manifold coordinates.
2.2 Multi-scale representations using SGW
In the last two decades, several multi-scale representations using irregular graphs have been suggested, including SGW [18] and Diffusion Wavelets [28]. In this work we focus on the multi-scale graph transform utilizing SGW. SGW provides a natural way of trading off between spectral and spatial resolution, since the SGW coefficients are localized in both domains. These wavelets are constructed using a kernel function $g(\mathbf{L})$ operator, which acts on a function by modulating each of its Fourier modes [18], that helps in capturing a trade off between vertex (spatial) and spectral localization. Spatial localization in the graph domain is implicitly controlled by a single wavelet scale parameter defined in the spectral domain, such that the more the vertex is localized in the vertex domain, the spectral band is wider. To represent a signal ${f}\in\mathbb{R}^{N}$ in multiple scales $S=[s_{1},s_{2},..s_{K}]$, the SGW transform is constructed as follows: Assume that $g(\lambda)$ is a filter in the spectral domain. Let $\delta_{i}\in\mathbb{R}^{N}$ be the delta function centred at the vertex $i\in G:\delta_{i}(j)=1\,\,\mbox{if}\,\,i=j$, and $\delta_{i}(j)=0$ otherwise. Given a bandpass filter $g(\lambda)$ and a wavelet centered on node $i$ at scale $s$, the wavelets $\psi(s,i),i=1,..N$ are calculated by applying them to the delta function on a single vertex $i$, given by $\psi(s,i)=\Phi g(s\lambda)\Phi^{T}\delta_{i}$. The value of $\psi(s,i)$ with respect to a vertex $m$ can be written as $\psi(s,i)(m)=\sum_{l=1}^{N}g(s\lambda_{l})\phi_{l}^{*}(i)\phi_{l}(m)$. Given a graph signal ${f}\in\mathbb{R}^{N}$, the SGW coefficient at node $i$ and scale $s$ can be expressed as follows:
$\psi_{f}(s,i)=\sum_{l=1}^{N}g(s\lambda_{l})\hat{f}(\lambda_{l})\phi_{l}(i)$ (1)
Fast computation using Chebyshev polynomials: Directly computing the SGW coefficients is computationally intensive - $O(N^{3})$ for $N$ points. Instead, Hammond et al. [18] suggested computing the SGW using a fast algorithm based on approximating the scaled generating kernels through low-order polynomials of $\mathbf{L}$ applied to the input data (details in supplementary material).
3 Our Proposed framework: Multi-Scale IMAP
In this section, we introduce Multi-scale IMAP, a framework for interpretable embedding via manifold learning utilizing a multi-scale graph representation. This approach enables us to maintain global regularity and preserve local structure without sacrificing interpretability. Our method imposes a differentiable structure on the mapping $h:\mathbf{x}\rightarrow\psi_{\mathbf{x}}$, supported on a discrete graph $G=(V,\mathbf{W})$, where $\psi_{\mathbf{x}}$ represents the encoded multi-scale graph transform of $\mathbf{x}$. Multi-scale IMAP consists of two main components:
Step 1: Features representation encoding using multi-scale graph representation.
In this step, our approach constructs a multi-scale graph representation by incorporating feature signals through the SGW Transform across multiple scales and graph frequencies. We suggest a 3D tensor-based approach (detailed in Algorithm 1), where we construct a 3D tensor with dimensions ${K}\times{N}\times{D}$ to provide an effective representation for encoding multi-scale transform. We leverage the encoded 3D tensor structure to align the optimized embedding with all scales simultaneously to enforce differentiable structure of the transformed features. Note that an alternative would be to use simple concatenation, all of the features transformed for all scales are represented by concatenating all the SGW coefficients for each point in a vector form. For completeness we also describe this alternative in the Appendix. The 3D tensor-based approach (detailed in Algorithm 1) provides the most effective representation for the following reasons:
(i) Joint and simultaneous optimization: using the 3D tensor-based method enables the joint and simultaneous optimization of all scales for each sample by leveraging the SGW coefficients of each spectral band within the cross-entropy loss framework while adhering to the fixed graph structure (Equation (3.2)). This approach effectively balances coarse and fine-grain resolutions, with summation performed only after the contrastive learning optimization process is complete. By smoothing both low- and high-frequency information jointly and simultaneously with respect to the graph structure, this method reduces the reliance on threshold-based techniques for removing high-frequency spectral bands, which may be less efficient. Jointly optimizing across scales ensures consistency in the optimization process for all scales with respect to a fixed pair of nodes. Although there may be correlations between overlapping SGWs corresponding to different spectral bands, each scale is associated with a distinct bandwidth of frequencies, effectively capturing graph structures at varying levels of neighborhood information. By integrating multiple scales, SGWs capture neighborhood information continuously across varying granularities, enabling a representation that strikes a balance between local and global structural details.
(ii) Enhanced Interpretability: the proposed representation establishes a one-to-one correspondence between input features and the final embedding features, making feature interpretation more straightforward and intuitive. This is demonstrated in Section 6, where we show that the embeddings produced enable effective feature importance rankings. By ensuring this direct correspondence, our approach facilitates meaningful interpretation of the embedding space. This combination of optimization efficiency and enhanced interpretability highlights the clear advantages of the 3D tensor-based representation.
Step 2: Optimization design using multi-scale network structure.
We integrate the multi-scale representation within the SGD optimization framework, simultaneously leveraging both low and high-frequency information. This integration leads to fine-grained manifold regularization and improved robustness in downstream tasks.
3.1 Step 1: Feature Representation to encode multi-scale structure
To encode multi-scale representations used for subsequent optimization we focus on a 3D tensor based method. Note that each dimension in the embedding space is constructed using a single feature in the original feature set, which is an essential characteristic that can be leveraged for interpreting graph embeddings.
3D-Tensor Encoding: The 3D-Tensor Encoding method concatenates the multi-scale representation for all features at a fixed scale, generating a matrix representation $\psi_{\mathbf{x}}(s_{j},:,:)\in\mathbb{R}^{D}\times\mathbb{R}^{N}$ for each scale $s_{j}$. After concatenating all $f_{i}$ at a fixed scale $s_{j}$, we have
$\psi_{\mathbf{x}}(s_{j},:,:)=\psi_{{f}_{1}}(s_{j},:)\|\psi_{{f}_{2}}(s_{j},:)\|...\|\psi_{{f}_{D}}(s_{j},:)$
[…]
$\mathbf{c}(\psi_{{f}_{i}}(s_{k},:,:),\psi_{{f}_{j}}(s_{k},:,:))=\psi_{{f}_{i}}(s_{k},:,:)\left|\right|\psi_{{f}_{j}}(s_{k},:,:)$
The optimization can be performed separately for each scale $s_{j}$, or jointly for all scales $s_{j}$ using the 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$ We detail the proposed feature representation encoding methods in the pseudo code algorithm 1.
  Input: Set of points $\left\{\mathbf{x}_{i}\right\}_{i=1}^{N}$.
  Output: Initial Node Embeddings ${\psi}_{\mathbf{x}}$
  Step 1: Construct $G=(V,\mathbf{W})$ from $\left\{\mathbf{x}_{i}\right\}_{i=1}^{N}$.
  Step 2: Construct the Laplacian $\mathbf{L}$.
  Step 3: Compute $\lambda_{\mbox{max}}(\mathbf{L})$.
1  Step 4: Initial Embedding construction:
2For $r=1,..,D$ associated with the feature signals $\left\{f_{r}\right\}$ do:
3Compute $\psi_{{f}_{r}}(s_{j},:)$ at scales $s_{j},j=1,...K$ using Chebyshev approximation.
  Step 5: Concatenate $\psi_{{f}_{r}}(s_{j},:)$ for $r=1,...,D$ in a matrix form $\psi_{\mathbf{x}}(s_{j},:,:)$ for each fixed scale $s_{j}$, representing all scales using 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
Algorithm 1 Encoding multi-scale structure in the optimized embeddings
3.2 Step 2: Optimization design using multi-scale network structure.
We propose to use optimization based on SGD that begins with the initial embedding within the spectral graph domain. While employing positive and negative sampling, our aim is to optimize the embedding space directly in the SGW domain, revealing the intrinsic structure of manifold data while retaining high-frequency information associated with the graph Laplacian. This optimization within the SGW domain provides more effective regularization, allowing for the direct removal of noise from each spectral band. We outline the steps of the proposed approach.
3D-Tensor Based Optimization for graph Embedding
The encoded manifold representation is a 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$. An SGD update rule at iteration $t$, applied to the 3D tensor, is:
$(\tilde{\psi}_{\mathbf{x}}^{(t+1)})_{i,j,k}=(\tilde{\psi}_{\mathbf{x}}^{(t)})_{i,j,k}-\alpha\frac{\partial\mathcal{L}}{\partial(\psi_{\mathbf{x}})_{i,j,k}}$ (2)
where $\alpha$ is the learning parameter. In this case, we employ the cross-entropy loss function:
$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})=\sum_{i,j,k}$ $\left(w_{ij}\log\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right.$
$\left.+(1-w_{ij})\log\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right)$ (3)
where $v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}=\frac{1}{1+||\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:)||^{2}}$. We apply the optimization based on SGD with respect to each scale $s_{k}$, where the gradient of the loss is approximated by:
$\frac{\partial\mathcal{L}\left(\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)\mid\mathbf{W}\right)}{\partial\psi_{\mathbf{x}_{i}}(s_{k},:,:)}$ $=\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}\,r_{ij}(s_{k})$ (4)
$\quad-\sum_{j}\frac{1}{\|r_{ij}(s_{k})\|^{2}}v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}\,r_{ij}(s_{k})$ (5)
where $r_{ij}(s_{k})=\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:)$.
The optimization process uses SGD with positive and negative sampling for all scales $s_{k}$ given a pair of positive or negative pair of nodes $i,j$.
We compute the final embedding by summing up all the optimized embeddings with respect to each scale $s_{k}$
$\tilde{\psi}_{\mathbf{x}}=\sum_{k}\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)$ (6)
Note that the dimensionality of the final embedding
$\tilde{\psi}_{\mathbf{x}}\in\mathbb{R}^{D}\times\mathbb{R}^{N}$, maintaining the same dimensionality as the original input, and thus there exists a one-to-one correspondence between the coordinates of the input features $\mathbf{x}$ and $\tilde{\psi}_{\mathbf{x}}$. This correspondence ensures that each dimension of the feature space is directly mapped to a single dimension in the embedding space $\tilde{\psi}_{\mathbf{x}}$.
Our optimization process involves using SGD with positive and negative sampling, similar to recent graph embeddings methods [33, 25]. Positive edge samples are associated with attraction (first term on the right hand side), while negative samples (second term on the right hand side) refer to a pair of nodes that are not connected in the graph, which create repulsion among dissimilar points.
The optimization can be performed jointly for all scales using the 3D tensor. Specifically, our implementation employs an optimization process that operates jointly across all scales, as described in the SGD update rule for iteration $t$ in Eq. (2). At each iteration $t$, associated with either a positive sample (an edge) or a negative sample (a pair of unconnected nodes), the update rule from Eq.(2). is applied. The derivative of the loss function, as detailed in Eq. (4), is computed for all scales $(s_{1},...s_{K})$, with the sign in Eq.(2) depending on whether the sample is positive or negative. Once the optimization concludes, the final embedding is derived from the 3D tensor by summing over the scales, as specified in Eq.(6). Encoding the representation in a 3D tensor offers two key advantages. First, is the computational efficiency: the update rule at each iteration, associated with a pair of nodes, is performed jointly and simultaneously for all scales. This avoids the need for independent optimization per scale, which would significantly increase computational complexity. The second is optimization alignment: jointly optimizing across scales ensures alignment in the optimization process for all scales with respect to a fixed pair of nodes (positive or negative sample). This approach is more effective than optimizing scales independently, as it maintains consistency and improves overall performance.
4 Theoretical results: Sampling set for Smooth Manifolds with functions defined over Paley-Wiener Spaces
In this section, we characterize the theoretical properties of the representation power of the SGW operator by considering functions sampled from the Paley-Wiener spaces [27] on combinatorial graphs. Pesenson introduced the Paley-Wiener spaces and demonstrated that Paley-Wiener functions of low type are uniquely determined by their values on certain subsets, known as uniqueness sets $U$. We show that the SGW operator can represent functions $f$ within the Paley-Wiener space more efficiently than the graph Laplacian operator $\mathbf{L}$. This efficiency is demonstrated by showing that the SGW operator is more effective in representing functions with larger bandwidth $\omega$ in the Paley-Wiener spaces (i.e., with higher frequencies) using subsets of nodes from the uniqueness sets $U$.
To characterize the representation properties of functions defined over $PW_{\omega}(G)$, we employ the notion of the $\Lambda$-set, introduced by Pesenson which is central to our investigation. Formally, the Paley-Wiener space of $\omega$-bandlimited signals is defined as $PW_{\omega}(G)=\left\{f\mid\hat{f}(\lambda)=0\,\,\forall\,\,\lambda>\omega\right\}$. The space $L_{2}(G)$ is defined as the Hilbert space of all complex-valued functions, and $L_{2}(S)$ is defined as the space of all functions from $L_{2}(G)$ with support in $S$: $L_{2}(S)=\left\{\varphi\in L_{2}(G)\mid\varphi(v)=0,v\in V(G)\setminus S\right\}$.
The $\Lambda$-set is defined as follows: a set of vertices $S\subset V$ is a $\Lambda(S)$-set if any $\varphi\in L_{2}(S)$ satisfies the inequality $||\varphi||\leq\Lambda||\mathbf{L}\varphi||$ for some constant $\Lambda(S)>0$. The infimum of all $\Lambda>0$ for which $S$ is a $\Lambda$-set is denoted by $\Lambda$.
The ability of the SGW operator to efficiently represent functions $f\in PW_{\omega}(G)$ can be summarized in the following theorem, which highlights the role of the $\Lambda_{\psi}$-set with respect to the operator $\psi$. We show that any signal $f\in PW_{\omega}(G)$, where $\lambda_{1}\leq\omega<\Omega_{G}$ for some $\Omega_{G}<\lambda_{N}$, can be uniquely represented by its samples on the uniqueness set $U$ using the SGW operator. Under certain conditions related to the SGW operator, its associated $\Lambda_{\psi}$-set is smaller than the $\Lambda$-set associated with the Laplacian operator.
Theorem 1
Let $G=(V,\mathbf{W})$ be a connected graph with $n$ vertices and $f\in PW_{\omega}(G)$ for $\lambda_{1}\leq\omega<\lambda_{\text{max}}$. The SGW operator $\psi$ can be constructed such that the set $S$ is a $\Lambda_{\psi}$-set and the set $U=V\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<1/\Lambda_{\psi}$ and $\Lambda_{\psi}<\Lambda$ for any $\varphi\in L_{2}(S)$, where $\Lambda$ is the $\Lambda$-set of the Laplacian operator.
The proof is provided in the Appendix.
Figure 1: Illustration of the proposed method. (Left) Initial encoding using SGW in (b) and regularized embedding in (c). (Right) Feature embedding space representation, which has a one-to-one correspondence with the input features, can be used along with clustering for interpretability models, such as feature importance.
5 Experimental Results
We demonstrate the advantage of our approach over existing approaches using both synthetic and real datasets. We evaluate our method by testing its clustering performance using the output embedding and comparing it to several representative methods, including UMAP, t-SNE, Diffusion Maps, ISOMAP, PHATE [26], PaCMAP [36], TriMap [1] and HeatGeo [22]. For UMAP, t-SNE, PHATE, PaCMAP, TriMap and HeatGeo, we hyperparameter tune the performance on each dataset, running the method five times for each hyperparameter configuration and taking the average. Then we report the best average score. Additional details are provided in the supplementary material.
Evaluation metrics: We test performance using the Adjusted Rand Index (ARI) and Adjusted Mutual Information (AMI). In all experiments we used $k$-means to cluster the data in the embedding space.
5.1 Real Datasets
We study the performance of MS-IMAP compared to other methods for real datasets. We chose a mix of datasets from varied fields: the Census dataset [14] is a financial dataset containing information about individuals extracted from the 1994 US Census; the Zilionis dataset [40] is a biological dataset containing single-cell sequencing data from different types of cells; the Animals with Attributes (AWA) [37] dataset is an image dataset containing images of animals; The Human Activity Recognition (HAR) [2] dataset which consists of sensor data collected from subjects performing six different activities: walking, walking upstairs, walking downstairs, sitting, standing, and lying. Each activity is represented as a feature vector of 561 attributes and 7352 samples; and the Cora dataset [30, 24] is a graph dataset of a citation network consisting of scientific publications. Note that we only use the features without the natural graph associated with Cora. More information about each dataset can be found in the supplementary material. As shown in Table 1, MS-IMAP is competitive and performs the best overall, while also providing an interpretable embedding through its one-to-one correspondence with the original features—unlike the other competing graph embedding methods.
Runtime and Computational Complexity The computational complexity of Multi-Scale IMAP is of $O(NDlog(N))$. The execution time of our method with Python code implementation with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM on the Cancer QC data-set of 48,969 samples and 306 features took 27 mins, on Cora 5 min, on the AWA dataset 1.35 mins and on Census 7 min. The increase in running times compared to methods such as UMAP is primarily due to our extension of manifold learning to include interpretability. Our approach generally requires the embedding dimensionality to be at least equal to the input dimensionality of the features, which adds computational complexity but ensures meaningful feature representation and interpretation. Our method is practical for handling datasets with millions of points and dozens to hundreds of features within a few hours. In fields such as finance and genomics/biology, where typical dataset sizes range from under 100,000 to several million data points, analysis is often conducted offline once the graph embedding is computed. Runtimes of up to a few hours are generally acceptable and do not present a significant challenge, making the method well-suited for real-world applications. Additional details are provided in the Appendix.
Data HAR Census Zilionis AWA Cora
Method / Accuracy ARI AMI ARI AMI ARI AMI ARI AMI ARI AMI
Features 0.45 0.58 0.11 0.15 0.48 0.63 0.07 0.16 0.01 0.07
ISOMAP 0.55 0.65 0.18 0.09 0.44 0.56 0.52 0.60 0.12 0.17
Diffusion Maps N/A N/A 0.06 0.16 0.01 0.02 0.22 0.42 0.03 0.07
UMAP 0.61 0.73 0.23 0.15 0.52 0.74 0.74 0.82 0.33 0.36
t-SNE 0.71 0.77 0.17 0.17 0.41 0.70 0.73 0.80 0.21 0.27
PHATE 0.55 0.66 0.20 0.09 0.65 0.72 0.66 0.73 0.12 0.22
TriMap 0.59 0.72 0.2 0.14 0.62 0.77 0.74 0.81 0.15 0.2
PaCMAP 0.62 0.75 0.24 0.15 0.55 0.76 0.75 0.81 0.21 0.26
HeatGeo 0.6 0.69 0.15 0.10 N/A N/A 0.65 0.74 0.17 0.20
MS-IMAP 0.67 0.8 0.23 0.15 0.70 0.76 0.74 0.81 0.33 0.36
Table 1: Clustering results comparison using ARI and AMI on the Census, Zilionis, AWA, and Cora datasets. Note HeatGeo had code execution issues on the Zilionis dataset.
6 Feature Importance using Explicit Correspondence with Embedding Features
The correspondence established between the features and the embedding space offers a valuable tool for assessing feature importance. We explore two alternative methods for measuring feature importance, leveraging the constructed embedding space, where each dimension can be traced back to its corresponding feature in the original space. Specifically, we test these methods in an unsupervised or self-supervised setting where the subset of features are selected without using label information.
6.1 Assessing Embedding Feature Importance using the Laplacian Score
The first method utilizes the Laplacian Score (LS) [19], which assesses the importance of each feature by examining its correlation with the eigenvalues of the graph Laplacian. The significance of the variable corresponding to the $l$-th coordinate, represented by the feature $f_{l}\in\mathbb{R}^{N}$, is inferred from the relevance of its corresponding coordinate $l$ in the embedding space, $\tilde{\psi}_{\mathbf{x}_{l}}$. The Laplacian Score is calculated with respect to the embedding features $\tilde{\psi}_{\mathbf{x}}$ using the graph Laplacian $\mathbf{L}$ and the degree matrix $\mathbf{D}$. To compute the Laplacian Score, we first subtract the mean and then calculate it as follows: $L_{s}(\tilde{\psi}_{\mathbf{x}})_{l}=\frac{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{L}(\tilde{\psi}_{\mathbf{x}})_{l}}{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{D}(\tilde{\psi}_{\mathbf{x}})_{l}}$
Smaller scores indicate that the embedding feature $(\tilde{\psi}_{\mathbf{x}})_{l}$ has a greater projection onto the subspace of eigenvectors corresponding to the smallest eigenvalues, signifying higher importance concerning the global graph structure. Therefore, the feature importance of $(\tilde{\psi}_{\mathbf{x}})_{l}$ can be directly interpreted as the importance of the corresponding original feature.
6.2 Assessing Importance of Embedding Features using Mutual Information
In the second method, we propose a self-supervised approach that uses Mutual Information (MI) to estimate the importance of each dimension in the embedding space, which can then be explicitly linked to the original features. First, we apply $k$-means to cluster the MS-IMAP embedding space and use the resulting clusters as pseudo-labels. We then compute MI to quantify the shared information between features and these pseudo-labels, similar to how MI would be computed with ground truth labels. This step allows MI to measure how much each feature contributes to the cluster assignments, highlighting those most informative in distinguishing clusters. The MI-Based Feature Importance identify features that are informative about cluster assignments. Thus, masking these features would especially degrade the clustering performance. We summarize the MI-based unsupervised feature selection in Algorithm 2. In comparison to the MI based feature importance method, using the Laplacian score [19] we have that the embedding may have combined original features in a way that makes individual Laplacian Scores less indicative of feature importance. The Laplacian score prioritize neighborhood preservation, which might not directly translate to cluster discrimination efficacy. On the other hand, while the proposed MI based feature importance identifies features that are directly informative for distinguishing clusters, one of its possible limitations may be relaying on the quality of pseudo-labels, which may not align with true clusters.
  Input: MS-IMAP embeddings $\psi_{\mathbf{x}}$, number of clusters $k$
1  Output: MI scores for each feature
  Step 1: Apply k-means clustering on the embedding space $\psi_{\mathbf{x}}$ to obtain cluster labels
[…]
  Step 2: Use the resulting cluster labels as pseudo-labels
[…]
  Step 3: Initialize an empty list to store mutual information (MI) values for each feature
[…]
  Step 4: For each feature $i$ in the original data or embedding space:
            Step 4.1: Compute mutual information between feature $\psi_{\mathbf{x}}(i)$ and the pseudo-labels
4            Step 4.2: $MI\_score=\text{MI}(\psi_{\mathbf{x}}(i),c({\psi_{\mathbf{x}}})$)
5            Step 4.3: Append $MI\_score$ to mi_scores
  Step 5: Output the list of MI scores
            Return mi_scores
Algorithm 2 MI-based Unsupervised Feature Selection
Figure 2: Clustering results on Cora using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
6.3 Assessing Feature Importance via Embedding Correspondence
We demonstrate the application of MS-IMAP for feature importance by evaluating clustering performance using $k$-means as a function of the number of features, based on the most significant features identified by two methods: the Laplacian score and Mutual Information-based feature selection. We compared the clustering results (using ARI and AMI) when $k$-means is applied to different subsets of embedding features selected by the Laplacian score and MI-based methods, with results obtained by applying the same feature selection methods directly to the original feature space.
Unlike most manifold learning and graph embedding methods, MS-IMAP produces embeddings where the dimensions can be directly mapped to the original feature space, enabling the assessment of individual feature contributions.
On the Cora dataset, Figure 2 shows that for both feature selection methods, the clustering accuracy obtained using the MS-IMAP embedding is significantly higher than when using the raw features across different subsets of features. This improvement is due to the robustness of the MS-IMAP embedding, while still offering interpretability by linking the embedding dimensions to the original features.
7 Discussion
Identifying the key factors of high-dimensional datasets is a necessary ingredient to take advantage of unlabelled data in many practical applications. In this article, we introduce a novel contrastive learning framework for manifold learning via graph embeddings, capitalizing on both low and high-frequency information. In particular, we use SGW to construct a multi-scale graph representation of the underlying input feature space. We then use an SGD-based optimization scheme together with innovative strategies such as 3D Tensor encoding to derive the embeddings. To motivate the use of the spectral graph wavelet (SGW) for representation, we examine the theoretical properties of the SGW by studying functions in Paley-Wiener spaces on combinatorial graphs. We prove that the SGW operator provides a more effective representation through the concept of the $\Lambda$-set. Finally, we demonstrate that the embeddings are interpretable using two alternative methods for unsupervised and self-supervised feature importance, applied to both the embedding and original feature spaces. The construction of our methodology inherently links the original and transformed feature spaces, a capability that is notably absent in most existing non-linear manifold learning techniques, to the best of our knowledge.
Limitations and future work include the following. Firstly, we rely on the assumption that the features, and thus the similarity measure, adequately capture information to reveal geodesic distances between manifold points. Secondly, utilizing the current optimization framework with SGD may present difficulties when extending to out-of-sample problems. Furthermore, although our embedding feature correspondence proves valuable in assessing feature importance through a global method, there remains potential for enhancement through the incorporation of local methods.
Appendix
8 Encoding Method 1
[…]
$\mathbf{c}(\psi_{{f}_{i}}(s_{k},:),\psi_{{f}_{j}}(s_{k},:))=\psi_{{f}_{i}}(s_{k},:)\left|\right|\psi_{{f}_{j}}(s_{k},:)$
denoting the concatenation of the vectors corresponding to the multi-scale representation $\psi_{{f}_{i}}(s_{k},:)$ and $\psi_{{f}_{j}}(s_{k},:)$. For method 1 all features and all scales are concatenated together, the resulting matrix $\psi_{\mathbf{x}}$ can be represented as
$\psi_{\mathbf{x}}=\psi_{{f}_{1}}(s_{1},:)\|\psi_{{f}_{2}}(s_{1},:)\|....\|\psi_{{f}_{D-1}}(s_{K},:)\|\psi_{{f}_{D}}(s_{K},:)$
where $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$.
1. Optimize Embedding Method 1:
Given the encoded multi-scale representation $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$ perform optimization in the SGW domain, using the following fuzzy cross entropy loss function:
$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})=\sum_{i,j}\left(w_{ij}\mbox{log}\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}}}+(1-w_{ij})\mbox{log}\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}}}\right)$ (7)
where $v_{ij}^{\psi_{\mathbf{x}}}=\frac{1}{1+||\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}}||^{2}}$.
Dropping terms that do not depends on $\psi_{\mathbf{x}_{i}}$, the gradient of the loss is approximated by:
$\frac{\partial\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})}{\partial\psi_{\mathbf{x}_{i}}}\sim\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}}(r_{ij})-\sum_{j}\frac{1}{||r_{ij}||^{2}}v_{ij}^{\psi_{\mathbf{x}}}(\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}})$ (8)
where $r_{ij}=\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}}$, and $\psi_{\mathbf{x}_{i}}$ corresponds to the vector associated with node $i$ multi-scale representation concatenated for all the graph signals and all spectral bands. since for method 1 all features and all scales are concatenated together,
Figure 3: The encoding step of our framework is illustrated as a mapping between each coordinate of the input features and a corresponding dimension in the embedding space. This mapping is facilitated through our proposed approach, which aligns the input dimensions of the features with the latent embedding space. By establishing this correspondence, we acquire the capability to interpret each dimension in the embedding space. Such interpretation allows for various analyses, including providing feature importance relative to the latent space.
9 Additional experiments: Embedding Feature Importance using Mutual Information and Laplacian Score
We include additional experiments that demonstrate the application of MS-IMAP for feature importance by evaluating clustering performance using $k$-means as a function of the number of features (in addtion to the one found in the Section 6), based on the most significant features identified by the Laplacian score and Mutual Information-based feature selection. We compared the clustering results (using ARI and AMI) when $k$-means is applied to different subsets of embedding features selected by the Laplacian score and MI-based methods, with results obtained by applying the same feature selection methods directly to the original feature space.
On the Zillions dataset, Figures 5 and 6 shows that for both feature selection methods, the clustering accuracy obtained using the MS-IMAP embedding is significantly higher than when using the raw features across different subsets of features. This improvement is due to the robustness of the MS-IMAP embedding, while still offering interpretability by linking the embedding dimensions to the original features. We also show in Figure 4 the clustering accuracy obtained using the MS-IMAP copared to using the raw features using AMI, demonstrating similar observations regarding the advantages of MS-IMAP.
Figure 4: Clustering results (AMI) on Cora using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
Figure 5: Clustering results (ARI) on Zillions using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
Figure 6: Clustering results (AMI) on Zillions using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
10 Dataset Details
Two Moons. The two moons dataset depicts two interleaving half-circles. We sampled $N=600$ points and used a Gaussian noise level having standard deviation $0.12$. An example is show in Figure 7. More specifically, to produce these points we use Sci-kit Learn’s sklearn.datasets.make_moons function with $n\_samples=600$ and $noise=0.12$.
In Figure 7, we show example clusterings for each of the methods mentioned in the Experimental Results section of the main paper.
Census. From the UCI Machine Learning Repository [14], this dataset contains 14 features that are a mix of categorical, numerical, and binary. Such features include age, marital status, sex, etc. The goal is to predict whether a sample makes less than or equal to \$$\$50,000$, or strictly more. We use $32,561$ samples in our dataset.
Lung Cancer. The Zilionis dataset is widely used, and consists of single-cell RNA sequencing data. It has $306$ features, and $48,969$ samples. The data has 20 classes corresponding to cell type. More can be found at [40]
Animals with Attributes. The Animals with Attributes (AWA) dataset, contains 5,000 data points corresponding to 10 unseen classes, where the testing image features are obtained from the pre-trained ResNet architecture, with $D=2,048$ dimensions, and the semantic features are provided with $D=85$ dimensions. More information can be found in Section 4.1 of [37].
Cora. The Cora dataset has 2,708 scientific publications categorized into seven classes. The network has 5,429 links and each publication is described as a binary word vector indicating the presence or absence of a word. The dictionary consists of 1,433 words. In our experiments, we do not use the given graph, and only use the features. [30, 24].
HAR. The Human Activity Recognition (HAR) dataset consists of sensor data collected from 30 subjects performing six different activities: walking, walking upstairs, walking downstairs, sitting, standing, and lying. The data was recorded using accelerometer and gyroscope sensors embedded in a smartphone worn on the waist. Each activity is represented as a feature vector of 561 attributes and 7352 samples, extracted from the raw sensor signals through signal processing techniques. In our experiments, we do not use the temporal dependencies and only utilize the extracted features.
Figure 7: Clustering results using two moons
11 Experimental Results on two moons Synthetic Dataset
11.1 Synthetic dataset
Here we show further experiments on more data, where we assess the robustness of our method using the two moons dataset, which is a 2D manifold depicting two interleaving half-circles. We sampled $N=600$ points, and set the Gaussian noise standard deviation to $0.12$. While spectral based methods such as UMAP are effective under relatively “modest” noise levels, their performance deteriorates in the presence of larger amounts of noise. As shown in Table 1 under the Moons column, our approach is robust and correctly clusters most points despite the large noise level, and competitive with the competing methods.
Data Dense-sparse clusters
Method/Accuracy ARI AMI
Features 0.24 0.18
UMAP 0.54 0.51
t-SNE 0.42 0.35
ISOMAP 0.36 0.3
Diffusion Maps 0.25 0.19
PHATE 0.42 0.35
HeatGeo 0.54 0.52
MS-IMAP 0.89 0.87
Table 2: Comparison of clustering performance on the two moons dataset.
12 Hyperparameter tuning details
We tune UMAP on each dataset in the Experimental Results section, by tuning over the parameter space in Table 3.
Hyperparameter Set of values
n_neighbors 2, 10, 15, 20, 30, 50, 100
min_dist 0, 0.1, 0.5, 0.99
n_components 2, 3
Table 3: Space of parameters in which we tuned UMAP.
12.1 TSNE Hyperparameter Tuning Details
We tune TSNE on each dataset in the Experimental Results section, by tuning over the parameter space in Table 4.
Hyperparameter Set of values
perplexity 15, 30, 60
early_exaggeration 12, 24
n_components 2, 3
Table 4: Space of parameters in which we tuned TSNE.
12.2 PHATE Hyperparameter Tuning Details
We tune PHATE on each dataset in the Experimental Results section, by tuning over the parameter space in Table 5.
Hyperparameter Set of values
knn 3, 5, 10
decay (i.e. $\alpha$) 6, 10, 14, 18, 40
t auto, 50, 100, 150, 200
Table 5: Space of parameters in which we tuned PHATE.
12.3 HeatGeo Hyperparameter Tuning Details
We tune HeatGeo on each dataset in the Experimental Results section, by tuning over the parameter space in Table 6,
Hyperparameter Set of values
knn 5, 10, 15
lap_type normalized, combinatorial
harnack_regul 0, 0.5, 1
Table 6: Space of parameters in which we tuned HeatGeo.
13 Ablation Studies of Hyperparameters for MS-IMAP
Here we study what effect the hyperparameters – the number of nearest neighbors, and the number of filters – have on the clustering performance of MS-IMAP with method 2. In table 7, we demonstrate the effect of varying the number of neighbors, between 10, 15, and 20 neighbors. We see that the results are mostly the same, thus showing MS-IMAP is robust on real datasets.
Dataset Census Zilionis AwA
Number of Neighbors / Accuracy ARI AMI ARI AMI ARI AMI
10 0.22 0.15 0.71 0.78 0.73 0.80
15 0.22 0.15 0.70 0.76 0.74 0.81
20 0.22 0.15 0.70 0.77 0.71 0.79
Table 7: Ablation study on MS-IMAP with encoding method 2, varying the number of neighbors. The number of filters is kept at 5.
We also study the effect of varying the number filters. In table 8, we also see similar results when using different filters, showing the stability of MS-IMAP.
Dataset Census Zilionis AwA
Number of Filters / Accuracy ARI AMI ARI AMI ARI AMI
5 0.22 0.15 0.70 0.76 0.74 0.81
6 0.22 0.15 0.70 0.76 0.73 0.81
7 0.22 0.15 0.72 0.78 0.74 0.81
Table 8: Ablation study on MS-IMAP with encoding method 2, varying the number of filters. The number of neighbors is kept at 15.
14 Ablation Studies of Hyperparameters for t-SNE, Isomaps, and Diffusion Maps
We study the effect of varying the number of neighbors for the methods: t-SNE, Isomaps, and Diffusion Maps. For t-SNE this would be the perplexity hyperparameter, for Isomaps this would be the number of neighbors, and for diffusion maps, this would be the parameter that affects the width of the Gaussian kernel, i.e. $\exp(\cdot/\alpha)$.
In Table 9, we see choosing a smaller perplexity of 15 does worse than the perplexity of 30 and 60.
Dataset Census Zilionis AwA
Perplexity / Accuracy ARI AMI ARI AMI ARI AMI
15 0.03 0.04 0.37 0.67 0.69 0.76
30 0.15 0.15 0.38 0.68 0.73 0.80
60 0.17 0.17 0.38 0.69 0.73 0.80
Table 9: Ablation study on t-SNE, varying the perplexity.
In Table 10, we see a similar pattern to t-SNE, where choosing too low causes reduction in performance. But performance stabilizes around choosing the number of neighbors as 5-10+.
Dataset Census Zilionis AwA
Neighbors / Accuracy ARI AMI ARI AMI ARI AMI
2 -0.05 0.02 0.41 0.55 0.43 0.58
5 0.18 0.09 0.44 0.56 0.52 0.60
10 0.18 0.09 0.44 0.57 0.51 0.62
Table 10: Ablation study on Isomap, varying the number of neighbors.
In Table 11, we find the ablation study for Diffusion Maps. Diffusion Maps perform the worst, but as with t-SNE and Isomap, benefits from considering more neighbors.
Dataset Census Zilionis AwA
alpha / Accuracy ARI AMI ARI AMI ARI AMI
1 0 0 0 0.01 0.23 0.44
2 0 0 0 0.01 0.22 0.43
3 0.06 0.16 0 0.02 0.22 0.42
Table 11: Ablation study on Diffusion Map, varying alpha which affects the width of the kernel.
15 Fast computation using Chebyshev polynomials
We provide additional details regarding the fast computation of SGW coefficients [18]. Directly computing the SGW coefficients above requires calculating the entire eigensystem of the Laplacian, which is computationally intensive - $O(N^{3})$ for $N$ points. Instead, Hammond et al. [18] suggested computing the SGW using a fast algorithm based on approximating the scaled generating kernels through low-order polynomials. The wavelet coefficients at each scale are then computed as a polynomial of $\mathbf{L}$ applied to the input data, using approximating polynomials given by truncated Chebyshev polynomials.
The Chebyshev polynomials $T_{k}(y)$ are computed using the recursive relations: $T_{k}(y)=2yT_{k-1}(y)-T_{k-2}(y)$ for $k\geq 2$, where $T_{0}=1$ and $T_{1}=y$.
The SGW coefficients are then approximated using wavelet and scaling function coefficients as follows:
$\psi_{f}(s_{j},i)\sim\left(\frac{1}{2}c_{j,0}f+\sum_{k=1}^{K}c_{j,k}\bar{T}_{j,k}(\mathbf{L})f\right)_{i}$ (9)
where $c_{j,k},j>0$ are the Chebyshev coefficients and $\bar{T}_{j,k}$ are the shifted Chebyshev polynomials $\bar{T}_{k}(x)=T_{k}\left(\frac{x-a}{a}\right)$ for $x\in[0,\lambda_{\mbox{max}}]$, where $x=a(y+1)$, a = $\lambda_{\mbox{max}}$. The scaling function coefficients, which are corresponding to a low-pass filter operation, are approximated in a similar way using Chebyshev polynomials. Note that the scaling kernel function is a low pass filter $h$ satisfying $h(0)>0$ and $h(x)\rightarrow 0$ when $x\rightarrow\infty$. If the graph is sparse, we obtain a fast computation of the matrix-vector multiplication $\bar{T}_{j,k}(\mathbf{L})f$, where the computational complexity scales linearly with the number of points, resulting in a complexity of $O(N)$ for an input signal $f\in\mathbb{R}^{N}$. The SGWs efficiently map an input graph signal (a vector of dimension $N$) to $NK$ scaling and wavelet coefficients.
16 Runtime and Computational Complexity
We have performed experimental runtime studies on empirical datasets. Experiments were performed on Virtual Server containers with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM. On the AWA dataset of $5,685$ samples and $2,048$ features, the 3D based Tensor and the simple concatenation Method both run in 1.35 mins. Finally on the Cancer QC data-set of 48,969 samples and 306 features, Method 1 took 12.44 mins. The computational complexity of Multi-Scale UMAP is of $O(NDlog(N))$ for construction of the multi-scale representations which includes the $k$ nearest neighbor graph using k-d tree, the SGW transform which is $O(N)$ for each dimension of the manifold for sparse graphs. The optimization stage has a complexity which scale with the number of edges in the graph, which has a complexity of $O(kDN)$.
17 Theoretical results: Sampling set for Smooth Manifolds with functions defined over Paley-Wiener Spaces
In this work, we characterize the theoretical properties of the representation power of the SGW operator by considering functions sampled from the Paley-Wiener spaces [27] on combinatorial graphs. The Paley-Wiener spaces were introduced on combinatorial graphs in [27] and a corresponding sampling theory was developed which resembles the classical one. Pesenson proved in [27] that Paley-Wiener functions of low type are uniquely determined by their values on certain subgraphs (which are composed from a set of nodes known as the uniqueness sets) and can be reconstructed from such sets in a stable way. We demonstrate that the SGW operator can represent functions $f$ that reside in the Paley-Wiener space on combinatorial graphs more efficiently than the graph Laplacian operator $\mathcal{L}$. The effectiveness of the SGW operator representation in this case can be understood in several ways. In one way, by the ability of the SGW operator to accurately represent functions with larger bandwidth, i.e., $f\in PW_{\omega^{\prime}}(G)$ where $\omega<\omega^{\prime}$. We first summarize the main notions and definitions. The space $L_{2}(G)$ is the Hilbert space of all complex-valued functions $f:V\rightarrow\mathbb{C}$ with the following inner product $\left\langle f,g\right\rangle=\sum_{v\in G}f(v)\overline{g(v)}$ and the norm
$||f||=\left(\sum_{v\in V}{|f(v)|^{2}}\right)^{1/2}$ (10)
The Laplace (normalized) operator $\cal L$ is defined by the formula [27]:
$\mathcal{L}f(v)=\frac{1}{\sqrt{d(v)}}\sum_{u\sim v}\left(\frac{f(v)}{\sqrt{d(v)}}-\frac{f(u)}{\sqrt{d(u)}}\right),f\in L_{2}(G)$ (11)
In order to prove these results, we will first need the following definitions:
Definition 1
The Paley-Wiener space of $\omega$bandlimited signals $f\in L_{2}(G)$ is defined as follows:
$PW_{\omega}(G)=\left\{f|\hat{f}(\lambda)=0\,\,\forall\,\,\lambda>\omega)\right\}$ (12)
We consider a simple, undirected, unweighted, and connected graph $G=(V,\mathbf{W})$, where $V$ is its set of $N$ vertices and $\mathbf{W}$ is its set of edges. The degree of $v$ is number of vertices adjacent to a vertex $v$ is and is denoted by $d(v)$. We assume that degrees of all vertices are bounded by the maximum degree denoted as
$d(G)=\mbox{max}_{v\in V}d(v)$ (13)
The following definition [27] explains the uniqueness set.
Definition 2
A set of vertices $U\subset V$ is a uniqueness set for a space $PW_{\omega}(G)$ if for every two functions from $PW_{\omega}(G)$ that coincide on $U$, then they coincide on $V$.
Definition 3
For a subset $S\subset V$, denote $L_{2}(S)$ as the space of all functions from $L_{2}(G)$ with support in $S$:
$L_{2}(S)=\left\{\varphi\in L_{2}(G),\varphi(v)=0,v\in V(G)\setminus S\right\}$
Definition 4
[27] We say that a set of vertices $S\subset V$ is a $\Lambda$set if for any $\varphi\in L_{2}(S)$ it admits a Poincare inequality with a constant $\Lambda=\Lambda(S)>0$
$||\varphi||\leq\Lambda||\mathbf{\cal L}\varphi||,\,\varphi\in L_{2}(S)$ (14)
The infimum of all $\Lambda(S)>0$ for which $S$ is a $\Lambda$set will be called the Poincare constant of the set $S$ and denoted by $\Lambda$.
The definition above provides a tool to determine when bandlimited signals in Paley-Wiener spaces $PW_{\omega}(G)$ can be uniquely represented from their samples on a given set. The role of $\Lambda$-sets was explained and proved in the following Theorem by Pesenson [27], that shows that if $S\subset V$ then any signal $f\in PW_{\omega}(G)$ can be uniquely represented by its function values in the complement set $U=V(G)\setminus S$:
Theorem 2
[27] If $S\subset V$ is a $\Lambda$- set, then the set $U=V(G)\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<\frac{1}{\Lambda}$.
Remark: Note that non-trivial uniqueness sets can not exist for functions from any Paley-Wiener subspace $PW_{\omega}(G)$ with any $\lambda_{0}\leq\omega\leq\lambda_{N}$, but they can exist for some range $\lambda_{0}\leq\omega<\Omega$, as was shown in [27].
We state one of our main results, in which we employ the SGW operator $\psi$ to characterize the uniqueness set using the $\Lambda_{\psi}$-set, therefore extending the $\Lambda$-set concerning the graph Laplacian operator $\cal L$.
Theorem 3
Let $G=(V,\mathbf{W})$ be a connected graph with $N$ vertices. Assume that there exist a set of vertices $S\subset V$ for which the conditions (1)-(2) in Lemma 1 below hold true. Let $\psi$ be the SGW operator using a polynomial $p(\mathbf{\cal L})$ with the coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that $\psi_{f}=\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}f$. Then, for any $\varphi\in L_{2}({S})$, we have that the following inequality holds:
$||\varphi||\leq\Lambda_{\psi}||\psi_{{\varphi}}||$ (15)
and thus the set $S$ is a $\Lambda_{\psi}$-set for the operator $\psi$ with $\Lambda_{\psi}={\frac{1}{\sqrt{\sum_{k=0}^{K}\frac{a_{k}^{2}}{\Lambda^{2k}}}}}$.
We recall the following results from [27]. Note that [27] established the construction of a $\Lambda$-set by imposing specific assumptions on the sets $S$ and $U$. Our result Theorem 3 holds similar assumptions as in [27].
Lemma 1 (Lemma 3.6 of [27])
Given a connected graph $G=(V,\mathbf{W})$, a set of vertices $S\subset V$, its complement $U=V\setminus S$, for which the following conditions hold true
  1. 1.
    For every $s\in S$ there exists $u\in U$ that is a neighbor of $s$, i.e., $w(u,s)>0$.
  2. 2.
    For every $s\in S$ there exists at least one neighbor node $u\in U$ whose adjacency set intersects $S$ only over $s$.
Then there exist a set of vertices $S\subset V$ which is a $\Lambda$-set, with $\Lambda=d(G)$.
We require one more property of the Laplacian operation, Lemma 3 below. To this end, we use a lemma from [27],
Lemma 2 (Lemma 3.9 of [27])
If $S$ is a $\Lambda$-set, then for any $\varphi\in L_{2}(S)$ and all $t\geq 0$, $k=2^{l}$, $l=0,1,2,...$
$\frac{1}{\Lambda^{k}}\|\mathcal{L}^{t}\varphi\|\leq\|\mathcal{L}^{k+t}\varphi\|,\quad\varphi\in L_{2}(S)$
Now we are ready to prove a modification of Lemma 3.9 from [27]:
Lemma 3
Let $S$ be a $\Lambda$-set, $\varphi\in L_{2}(S)$, and let $m\geq 1$ be an integer. Then
$\frac{1}{\Lambda^{m}}\|\varphi\|\leq\|\mathcal{L}^{m}\varphi\|.$
Proof: Because every positive integer has binary representation, then we have for some $j_{i}\in\{0,1,2,...\},i=1,2,...,N$, that
$m=\sum_{i=1}^{N}2^{j_{i}}.$
Using Lemma 2, let $k=2^{j_{1}}$ and $t=2^{j_{2}}+\cdots+2^{j_{N}}$, so $m=k+t$. Then we have
$\|\mathcal{L}^{m}\varphi\|$ $=\|\mathcal{L}^{k+t}\varphi\|$
$\geq\frac{1}{\Lambda^{k}}\|\mathcal{L}^{t}\varphi\|$
$=\frac{1}{\Lambda^{2^{j_{1}}}}\|\mathcal{L}^{2^{j_{2}}+\cdots+2^{j_{N}}}\varphi\|$
We can now let $k_{2}=2^{j_{2}}$ and $t_{2}=2^{j_{3}}+\cdots+2^{j_{N}}$ so that,
$\|\mathcal{L}^{m}\varphi\|\geq\frac{1}{\Lambda^{2^{j_{1}}}}\|\mathcal{L}^{2^{j_{2}}+\cdots+2^{j_{N}}}\varphi\|$ $=\frac{1}{\Lambda^{2^{j_{1}}}}\|\mathcal{L}^{k_{2}+t_{2}}\varphi\|$
$\geq\frac{1}{\Lambda^{2^{j_{1}}}}\frac{1}{\Lambda^{k_{2}}}\|\mathcal{L}^{t_{2}}\varphi\|$
$=\frac{1}{\Lambda^{2^{j_{1}}+2^{j_{2}}}}\|\mathcal{L}^{2^{j_{3}}+\cdots+2^{j_{N}}}\varphi\|$
Continuing in this way, we eventually get,
$\|\mathcal{L}^{m}\varphi\|\geq\frac{1}{\Lambda^{m}}\|\varphi\|.$
completing the proof. $\square$
In the next theorem, we expand the characterization of the uniqueness set using the $\Lambda$-set concerning the graph Laplacian operator $\mathcal{L}$ to include cases where we employ the SGW operator $\psi$. We thereby characterize the uniqueness set using the $\Lambda_{\psi}$-set for the SGW operator.
We now turn to prove Theorem 3, which was stated earlier.
Proof of Theorem 3:
Let $S\subset V$ be a set satisfying the conditions in Lemma 1, so $S$ is $\Lambda$-set. Then by Lemma 3, for any integer $k\geq 1$,
$\frac{1}{\Lambda^{k}}\|\varphi\|\leq\|\mathcal{L}^{k}\varphi\|.$
The SGW operator of $\varphi$ has the form,
$\psi_{\varphi}:=\sum_{i,j}\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}\varphi$
for some $K\geq 1$, $a_{k}\in\mathbb{R}$ are coefficients, and where the summation over $i$ and $j$ represents the summation over the nodes of the graph. Now we calculate,
$\|\psi_{\varphi}\|$ $=\left(\sum_{i,j}\sum_{k=0}^{K}\left|a_{k}(\mathcal{L}^{k}\varphi)_{ij}\right|^{2}\right)^{1/2}$
$=\left(\sum_{i,j}\sum_{k=0}^{K}a_{k}^{2}\left|(\mathcal{L}^{k}\varphi)_{ij}\right|^{2}\right)^{1/2}$
$=\left(\sum_{k=0}^{K}a_{k}^{2}\sum_{i,j}\left|(\mathcal{L}^{k}\varphi)_{ij}\right|^{2}\right)^{1/2}$
$=\left(\sum_{k=0}^{K}a_{k}^{2}\|\mathcal{L}^{k}\varphi\|^{2}\right)^{1/2}$
$\geq\left(\sum_{k=0}^{K}a_{k}^{2}\frac{1}{\Lambda^{2k}}\|\varphi\|^{2}\right)^{1/2}$
$=\|\varphi\|\left(\sum_{k=0}^{K}a_{k}^{2}\frac{1}{\Lambda^{2k}}\right)^{1/2}$
Letting
$\Lambda_{\psi}:=\frac{1}{\sqrt{\sum_{k=0}^{K}\frac{a_{k}^{2}}{\Lambda^{2k}}}}$
Then we have $\|\varphi\|\leq\Lambda_{\psi}\|\psi_{\varphi}\|$, so $S$ is also a $\Lambda_{\psi}$-set.
$\square$
From the assumptions and proof of Theorem 3, we get the immediate result:
Corollary 1
Let $S\subset V$ be a $\Lambda$-set. Then $S$ is also a $\Lambda_{\psi}$-set.
Remark 1: An important property which can be observed from Theorem 3 is the following: given the SGW $\psi_{{f}}$ as a spectral representation operator applied on $f$, we can choose coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that we obtain a $\Lambda_{\psi}$ set associated with the operator $\psi$, which is smaller than the $\Lambda$ set associated with the Laplacian operator $\cal L$. This implies that the operator $\psi$ provides more flexibility and better control over smoothness properties in comparison to the Laplacian operator.
Remark 2: Since the $\Lambda_{\psi}$ set can be chosen to be smaller than $\Lambda$ set (for a proper choice of the coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ using the operator $\psi$) then the SGW operator $\psi$ provides a more efficient representation for $f\in PW_{\omega^{\prime}}(G)$ with $\omega<\omega^{\prime}$ using the same subsets of nodes from the uniqueness set $U$ in comparison to the Laplacian operator $\cal L$.
Remark 3: Note that the characterization of the the uniqueness set does not rely on a reconstruction method of the graph signal values of $f(S)$ from their known values on $U$.
The next Theorem demonstrates the role of the $\Lambda_{\psi}$set with respect to the operator $\psi$, where we show that any signal $f\in PW_{\omega}(G)$ can be uniquely represented by its samples on the uniqueness set $U$. This results resembles the role of $\Lambda$-sets with respect to the graph Laplacian operator $\mathbf{\cal L}$, yet with a different bound then Lemma 2.
Theorem 4
Let $G=(V,\mathbf{W})$ be a connected graph with $N$ vertices and $f\in PW_{\omega}(G)$ for $\lambda_{1}<\omega<\lambda_{\mbox{max}}$. Given the SGW operator $\psi$, and a set $S$ which is a $\Lambda_{\psi}$ set. Then the set $U=V\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<1/\Lambda_{\psi}$.
Proof: Given $f,\tilde{f}\in PW_{\omega}(G)$, then $f-\tilde{f}\in PW_{\omega}(G)$. Assume that $f\neq\tilde{f}$. If $f,\tilde{f}$ coincide on $U=V\setminus S$, then $f-\tilde{f}\in L_{2}(S)$ and therefore
$||f-\tilde{f}||\leq\Lambda_{\psi}||\psi_{f-\tilde{f}}||$ (16)
Since $\psi_{{f}}\in\mathbb{R}^{N}$, we have that by properties of a vector space in $\mathbb{R}^{N}$, using the Cauchy–Schwarz inequality and assuming $|a_{k}|\leq 1\,\forall k$, we have:
$||\psi_{f-\tilde{f}}||\leq\omega||f-\tilde{f}||$ (17)
Combining the inequalities above and using the inequality $\Lambda_{\psi}\omega<1$ we have that
$||f-\tilde{f}||\leq\Lambda_{\psi}||\psi_{f-\tilde{f}}||\leq\Lambda_{\psi}\omega||f-\tilde{f}||<||f-\tilde{f}||$ (18)
which is a contradiction to the assumption that $f\neq\tilde{f}$. Thus, the set $U=V\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<1/\Lambda_{\psi}$. $\square$
Remark 1: Note that $\Lambda\omega<1$ implies that $\Lambda_{\psi}\omega<1$ given $\Lambda_{\psi}<\Lambda$, then we can increase the size number of nodes in the uniqueness set $U$ for $PW_{\omega}(G)$ (for a proper choice of the coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ using the operator $\psi$). In other words, we may increase the size of $S$ (thus reducing the size of $U$) and still obtain a uniqueness set with a smaller size for the graph signals in $PW_{\omega}(G)$.
Remark 2: We note that the results of Theorem 4 concerning the uniqueness set are independent from the stability properties of the representation. In order to achieve stability which is important for reconstruction, it is required to construct a wavelet operator using multiple scales $t_{j},j=1,...J$, as proposed in [18]. More specifically, we can express the function $\varphi$ using multiple scales $t_{j}$ (here we replace the previous notation of scale $s_{j}$ with $t_{j},j=1,..J$, not to confuse with nodes $s\in S$). Then for a fixed scale $t_{j}$ we have that the SGW is given by $||\psi_{\varphi}(t_{j})||=\left(\sum_{i=1}^{N}\sum_{k=0}^{K}|a_{t_{j},k}({\cal L}^{k}\varphi(i))|^{2}\right)^{1/2}$ and
$||\psi_{\varphi}||_{=}\left(\sum_{j}\sum_{i=1}^{N}\sum_{k=0}^{K}|a_{t_{j},k}({\cal L}^{k}\varphi(i))|^{2}\right)^{1/2}$ (19)
In a similar way to the arguments provided in Theorem 3 we can choose coefficients $a_{t_{j},k}$ associated with the Laplacian polynomial such that the inequality $||\varphi||\leq\Lambda_{\psi}||\psi_{{\varphi}}||$ is satisfied.
Remark 3: Optimal Choice of coefficients $a_{k}$ for Minimizing $\Lambda_{\psi}$ for optimal choice of $\Lambda_{\psi}$ in the multi-scale setting.
To find the coefficients $a_{t_{j},k}$ that minimize $\Lambda_{\psi}$ where:
$\Lambda_{\psi}=\frac{1}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}}}$ (20)
subject to the constraint: $\sum_{j=1}^{J}\sum_{k=0}^{K}a_{t_{j},k}^{2}=1$, we optimize the coefficients $a_{t_{j},k}$, which control how much of the signal at scale $t_{j}$ and power $k$ is captured by the SGW operator. We use the Lagrangian $\mathcal{F}$, incorporating the energy constraint into the optimization problem and Define:
$\mathcal{F}(a_{t_{1},0},\dots,a_{t_{J},K},\mu)=\frac{1}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}}}+\mu\left(\sum_{j=1}^{J}\sum_{k=0}^{K}a_{t_{j},k}^{2}-1\right)$ (21)
[…]
Taking partial derivatives, solving for $a_{t_{j},k}$, and differentiating $\mathcal{L}$ with respect to each coefficient $a_{t_{j},k}$:
$\frac{\partial\mathcal{F}}{\partial a_{t_{j},k}}=\frac{-a_{t_{j},k}/\Lambda^{2k}}{\left(\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}\right)^{3/2}}+2\mu a_{t_{j},k}=0.$ (22)
rearranging the equation above we obtain:
$\frac{a_{t_{j},k}}{\Lambda^{2k}}=2\mu a_{t_{j},k}\left(\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}\right)^{3/2}$
assuming $a_{t_{j},k}\neq 0$, we can divide both sides of the equation by $a_{t_{j},k}$, yielding:
$\frac{1}{\Lambda^{2k}}=2\mu\left(\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}\right)^{3/2}$ (23)
thus the sum $\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}$ is proportional to the inverse of $\Lambda^{2k}$, implying that the optimal wavelet coefficients should decay proportionally to the inverse powers of the Laplacian eigenvalues $a_{t_{j},k}\propto\frac{1}{\Lambda^{k}}$.
Next, applying the energy constraint: $\sum_{j=1}^{J}\sum_{k=0}^{K}a_{t_{j},k}^{2}=1$ and substituting $a_{t_{j},k}=\frac{1}{\Lambda^{k}}$, we obtain:
$\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{1}{\Lambda^{2k}}=1$ (24)
Thus, the normalized optimal coefficients $a_{t_{j},k}$ are:
$a_{t_{j},k}=\frac{\frac{1}{\Lambda^{k}}}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{1}{\Lambda^{2k}}}}.$ (25)
Note that the wavelet energy is distributed across all scales and powers of the Laplacian, with a decay that prioritizes low-frequency components (small $k$).
Using the form of the coefficients $a_{t_{j},k}$ above, the expression for $\Lambda_{\psi}$ is:
$\Lambda_{\psi}=\frac{1}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{1}{\Lambda^{4k}}}}$ (26)
which is the minimized value of $\Lambda_{\psi}$, showing that the optimal choice of coefficients reduces the overall constant, improving the localization and efficiency of the SGW operator. Thus, the optimal choice of wavelet coefficients $a_{t_{j},k}$ that minimizes $\Lambda_{\psi}$ in the multi-scale SGW operator is proportional to $\frac{1}{\Lambda^{k}}$, prioritizing low-frequency components.